OpenStudy (anonymous):

What is the equation of the line passing through (–3, 6) and parallel to the line y = 4x – 2 in standard form? x – 4y = –18 x + 4y = 18 4x – y = –18 4x + y = 18

OpenStudy (zepp):

I need to post at least once before @waterineyes :(

OpenStudy (anonymous):

Parallel lines have equal slope.. y = 4x - 2 Slope = 4 So use previous equation now.. y - 6 = 4(x + 3) 4x - y = -18

OpenStudy (anonymous):

@zepp I did not get what you said..

OpenStudy (anonymous):